# Tokenization is Sensitive to Language Variation

### v1

@@ 8 Conclusion
In this work, we investigated tokenizer settings for tasks that require robustness and tasks that require sensitivity to language variation. BPE settings perform differently on the two task types. We make three practical suggestions for selecting tokenizer settings: (1) Pay the most attention to the pre-tokenizer. It influences how Unicode Character Categories can be combined (e.g., “b4” in Table 4), and how many different words are ultimately part of the vocabulary (e.g., “_queer” in Table 4). (2) Choose a bigger vocabulary size for settings that require sensitivity to language variation. (3) Use a small machine learning classifier to test the effect of different tokenizers on tasks in your domain of interest. In the future, it might be helpful to make intrinsic measures sensitive to tasks when used to estimate model performance. With this work, we hope to encourage further work on language variation and its relation to tokenizers to ultimately build better, fairer and more robust LLMs.
@@
We evaluate tokenizers on 110M parameter encoder models as encoder models tend to reach higher performance for classification tasks in low parameter settings. However, we risk not accounting for emergent capabilities of popular larger generative language models. For example, the number of context tokens for recent models are generally much larger than the original 512 tokens of BERT base.
We find that logistic regression correlates with downstream performance on classification tasks. For more complex tasks like regression or generation tasks, logistic regression might not be applicable. However, we still see it as a great benefit that logistic regression can enable us to quickly test tokenizers on a variety of different tasks and make a more informed decision on what tokenizers to test more rigorously.
For tasks requiring robustness to language variation, many more tasks could have been included. Future work could investigate tasks such as question answering or arithmetic tasks. For tasks requiring sensitivity to language variation, we originally planned to include more classification tasks that are tied to well-established factors of language variation, e.g., age prediction and geographic location prediction Nguyen et al. (2016). However, we repeatedly encountered difficulties with cleanly separating semantic from form-based cues. For age prediction and location prediction specifically, logistic regression models made extensive use of content words. Based on the coefficients of our logistic regression models, we excluded tasks that we expected to mainly require sensitivity to language variation, but where the models primarily relied on content words.
The tokenizer settings seem to have different effects when the pre-training corpus and fitting corpus differ. The pre-training corpus might influence in how far a model can leverage the diversity encoded in the tokenizer. See Appendix LABEL:app:webbook-pretraining for further experiments. Note, however, that recent work suggests to only use vocabulary that appears in the training corpus Land and Bartolo (2024) – which is unlikely when different corpora are used for model training and tokenizer fitting.
@@ Acknowledgements
We thank the colleagues in the various NLP Groups at Utrecht University and the Blablablab at the University of Michigan and, specifically, Kees van Deemter, Antal van den Bosch, Yupei Du, Qixiang Fang, Melody Sepahpour-Fard, Shane Kaszefski Yaschuk, Elize Herrewijnen, Sanne Hoeken, Hugh Mee Wong, Jian Zhu and Pablo Mosteiro for, among others, feedback on writing and presentation. We thank colleagues at EMNLP 2024 for their insights into tokenization, and, specifically, Vilém Zouhar. We thank Laura Wegmann for feedback on writing.

### v2

@@ 8 Conclusion
In this work, we investigated tokenizer settings for tasks that require robustness and tasks that require sensitivity to language variation. BPE settings perform differently on the two task types. We make three practical suggestions for selecting tokenizer settings: (1) Pay the most attention to the pre-tokenizer. It influences how Unicode character categories can be combined (e.g., “b4” in Table 4), and how many different words are ultimately part of the vocabulary (e.g., “_queer” in Table 4). (2) Choose a bigger vocabulary size for settings that require sensitivity to language variation. (3) Use a small machine learning classifier — e.g., a logistic regression classifier — to evaluate how different tokenizers affect performance on tasks in your domain. For example, for general-purpose language models, pick a tokenizer that this model consistently predicts to perform well across both variation-robust and variation-sensitive tasks. The best tokenizer settings might be sensitive to what types of language variation (e.g., lexical vs. syntactic) are present. We think it is crucial to investigate different types of language variation individually in future work to ultimately build better, fairer and more robust LLMs.
@@
We evaluate tokenizers on 110M parameter encoder models, because they tend to reach higher performance for classification tasks in low parameter settings. However, we risk not accounting for emergent capabilities of popular larger generative language models. For example, the number of context tokens for recent models are generally much larger than the original 512 tokens of BERT base.
We find that logistic regression correlates with downstream performance on classification tasks. For more complex tasks like text generation, logistic regression probably will not be applicable. However, we still see it as a great benefit that logistic regression can enable us to quickly test tokenizers on a variety of different tasks and make a more informed decision on what tokenizers to test more rigorously.
For tasks requiring robustness to language variation, many more tasks could have been included. For example, future work could include challenge NLI datasets, e.g., ANLI Nie et al. (2020). Further, future work could investigate different types of tasks such as question answering or arithmetic tasks. For tasks requiring sensitivity to language variation, we originally planned to include more classification tasks that are tied to well-established factors of language variation, e.g., age prediction and geographic location prediction Nguyen et al. (2016). However, we repeatedly encountered difficulties with cleanly separating semantic from form-based cues. For age prediction and location prediction specifically, logistic regression models made extensive use of content words. Based on the coefficients of our logistic regression models, we excluded tasks that we expected to mainly require sensitivity to language variation, but where the models often relied on content words.
We varied one tokenizer setting at a time while leaving the other two on their “default values” (i.e., miscellaneous fitting corpus, gpt2 pre-tokenizer and vocabulary size of 32k), in order to isolate the effect of each individual setting without exhaustive testing of all possible combinations. Future work could explore interdependencies between tokenizer settings thorugh more comprehensive testing.
The tokenizer settings seem to have different effects when the pre-training corpus and fitting corpus differ. Specifically, the pre-training corpus might influence in how far a model can leverage the diversity encoded in the tokenizer. See Appendix LABEL:app:webbook-pretraining for further experiments. Note, however, that recent work suggests to only use vocabulary that appears in the training corpus Land and Bartolo (2024) – which is unlikely when different corpora are used for model training and tokenizer fitting.
@@ Acknowledgements
We thank the anonymous ARR reviewers for their constructive comments. We thank the colleagues in the various NLP Groups at Utrecht University and the Blablablab at the University of Michigan and, specifically, Kees van Deemter, Antal van den Bosch, Yupei Du, Qixiang Fang, Melody Sepahpour-Fard, Shane Kaszefski Yaschuk, Elize Herrewijnen, Sanne Hoeken, Hugh Mee Wong, Jian Zhu and Pablo Mosteiro for, among others, feedback on writing and presentation. We thank colleagues at EMNLP 2024 for their insights into tokenization, and, specifically, Vilém Zouhar. We thank Laura Wegmann for feedback on writing. This research is supported by the “Digital Society - The Informed Citizen” research programme, which is (partly) financed by the Dutch Research Council (NWO), project 410.19.007. This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ A.1 Fitting Corpora
Source Train Dev Test
Wikipedia 1,469,999,792 15,000,029 15,000,087
Twitter 1,470,004,662 15,000,048 15,000,057
PubMed 1,469,999,499 15,000,106 15,000,501
Mixed 1,477,872,323 15,080,505 15,080,919
Table 7: Fitting corpora with similar word counts. We compare three fitting corpora for tokenizers. Word count is calculated using white-space splitting. The size of the fitting corpora are not exactly the same when it comes to word count. But variations in word count are below 1% and should not affect the vocabulary of the tokenizer fitted on them.
Source Text word count domain
Wiki Mary Jane Christie Serrano (c. 1840 – 1923) was a writer, poet and considered … 24 -
Twitter Where are the top places in Broward or Palm Beach? [thinking-face][eyes] 10 -
PMed … Myoelectrical activity of the gut has been studied in the postoperative period … 132 -
Misc. Israel, as usual, wants American forces to fight a bloody war against Iran. … 66 nytimes
>In Israel, my grandfather fought for its life. The people down the street fought … 581 reddit
[…]
I read the audio version of this story and loved it. … 53 goodreads
[…]
there is Sydney waiting to enter. Their eyes meet. “Maggie! I was just … 998 ao3
… Abstract. How many n-orthants can be intersected in the n-dimensional … 2048 s2orc
[…]
if you’re looking for enhancement cores in the game one really useful way … 102 YouTube
I have a lot of ties. But my favorite one. My favorite tie is owned by … 183 blogcorpus
Witnesses interviewed: 3 (N°1141, 1142, 1143). Nikolay S.( N°1142): "After the … 72 Pile-CC
It came in perfect condition and it is very soft. 9 amazon
I’m sure that the procurement people are doing the best job they can … 94 sfu-socc
… Sort -11, -1, 0, -3, 5 in decreasing order … 1645 DM Maths
… "But the killer isn’t the Russian army." "It’s the subzero temperatures." … 1978 OpenSubtitles
[…]
Rumson, NJ – December 2013 What started in 2003 as a group of mostly Christian … 592 realnews
Table 8: Dataset examples. We show text examples for all used fitting corpora.
Genre Domain Train Dev Test
Forum Reddit 245M 2.5M 2.5M
Literature AO3 147M 1.5M 1.5M
Literature Gutenberg before 1919 49M 0.5M 0.5M
News Realnews 147M 1.5M 1.5M
News/Comments NYTimes & Comments 24M 0.3M 0.3M
News/Comments SFU-SOCC 3M 0.03M 0.02M
Q&A StackExchange 196M 2.0M 2.0M
Reviews Goodreads 49M 0.5M 0.5M
Reviews Amazon 49M 0.5M 0.5M
Mails Gmane 147M 1.5M 1.5M
Transcripts YouTubeCommons 98M 0.9M 1.0M
Transcripts OpenSubtitles 49M 0.5M 0.5M
Code GitHub 49M 0.5M 0.5M
Science S2ORC 98M 1.0M 0.9M
Blogs BlogCorpus 10M 0.1M 0.1M
Raw Text Webpages CommonCrawl 98M 1.0M 1.0M
Mathematics DM Mathematics 20M 0.2M 0.2M
Total: 1,478M 15.1M 15.1M
Table 9: Miscellaneous dataset statistics.